Green Mining Company has purchased a tract of mineral land for $954,000. It is estimated that this tract will yield 127,200 tons of ore with sufficient mineral content to make mining and processing profitable. It is further estimated that 6,360 tons of ore will be mined the first and last year and 12,720 tons every year in between. (Assume 11 years of mining operations.) The land will have a salvage value of $31,800.
The company builds necessary structures and sheds on the site at a cost of $38,160. It is estimated that these structures can serve 15 years but, because they must be dismantled if they are to be moved, they have no salvage value. The company does not intend to use the buildings elsewhere. Mining machinery installed at the mine was purchased secondhand at a cost of $63,600. This machinery cost the former owner $159,000 and was 50% depreciated when purchased. Green Mining estimates that about half of this machinery will still be useful when the present mineral resources have been exhausted, but that dismantling and removal costs will just about offset its value at that time. The company does not intend to use the machinery elsewhere. The remaining machinery will last until about one-half the present estimated mineral ore has been removed and will then be worthless. Cost is to be allocated equally between these two classes of machinery.

As chief accountant for the company, you are to prepare a schedule showing estimated depletion and depreciation costs for each year of the expected life of the mine. (Round per unit answers to 2 decimal places, e.g. 0.45 for computational purposes and final answers to 0 decimal places, e.g. 45,892.)

Estimated depletion cost

Year

Depletion

1st Yr. $

2nd Yr.

3rd Yr.

4th Yr.

5th Yr.

6th Yr.

7th Yr.

8th Yr.

9th Yr.

10th Yr.

11th Yr.

Estimated depreciation cost

Year

Building

Machinery (1/2)

Machinery (1/2)

1st Yr. $

$

$

2nd Yr.

3rd Yr.

4th Yr.

5th Yr.

6th Yr.

7th Yr.

8th Yr.

9th Yr.

10th Yr.

11th Yr.

ALSO compute the depreciation and depletion for the first year assuming actual production of 5,300 tons. Nothing occurred during the year to cause the company engineers to change their estimates of either the mineral resources or the life of the structures and equipment. (Round per unit answers to 2 decimal places, e.g. 0.45 for computational purposes and final answers to 0 decimal places, e.g. 45,892.)
Depletion $

?????????

Depreciation $

?????????
Answer #1

1. Depreciation per ton = ( 954,000 - 31800) / 127200 = 7.25 per ton

Estimated depletion cost
Year Depletion
1 46,110 [6360*7.25]
2 92220 [12720*7.25]
3 92220
4 92220
5 92220
6 92220
7 92220
8 92220
9 92220
10 92220
11 46,110 [6360*7.25]

2. Depreciation per ton for Building = 38160 /127200 = 0.3 per ton [ 1st & 11th year =6360*0.3 =1908 ; 2- 10 = 12720*0.3 = 3816]

Machinery ( 1/2) depreciation per ton =31800 / 127200 = 0.25 per ton [ 1st & 11th year =6360*0.25 =1590 ; 2- 10 = 12720*0.25 = 3180]

Machinery ( 1/2) depreciation per ton =(31800 / 127200) *2 = 0.50 per ton [ 1st & 6th year =6360*0.25 *2 =3180 ; 2- 5 = 12720*0.25*2 = 6360]

Building Machinery (1/2) Machinery (1/2)
1 1908 1590 3180
2 3816 3180 6360
3 3816 3180 6360
4 3816 3180 6360
5 3816 3180 6360
6 3816 3180 3180
7 3816 3180 0
8 3816 3180 0
9 3816 3180 0
10 3816 3180 0
11 1908 1590 0

3. Depletion = 7.25 * 5300 = 38425

Depreciation = 5300 * 0. 3+ 5300 *0.25 + 5300 * 0.25* 2 = 5565